OpenStudy (anonymous):

How do I Find the arclength of the curve r(t) = < 6sqrt(2t), e^(6t), e^(-6t)>. 0 =< t =< 1

OpenStudy (anonymous):

I believe you have to do this integral \[\int\limits_{0}^{1} \sqrt{72t+e^{12t}-e^{-12t}} \] There's no exact solution so you have to use simpson's rule.

OpenStudy (anonymous):

so i did find that integral and tried to solve it but my final result was wrong since the system say its wrong

OpenStudy (anonymous):

hey you guys kno da example of a reciprocal

OpenStudy (anonymous):

Ooops I made a mistake. Here's the general formula. \[\int\limits_{}^{} \sqrt{(dx/dt)^2+(dy/dt)^2+(dz/dt)^2}\]